# Autoregressive models for panel data causal inference with application to state-level opioid policies

### v1

@@ 5 Simulation studies
We evaluate the performance of the proposed approach against state-of-the-art methods for panel data causal inference in a simulation study. Rather than simulating outcomes from a data generating process we choose, we instead utilize existing outcome data from our application to ensure that the results are more relevant to our setting. Specifically, we let our observed outcome data represent $Y_{it}(0)$, and then we simulate treatment effects to obtain $Y_{it}(1)$. Our data consist of 51 units and 16 time periods, and we choose 20 units to be treated in all simulations. Our observed outcomes are the number of opioids prescribed by each state (and Washington DC) and year. We do not vary the outcome values across simulations, but instead vary which units are treated, when they are treated, and the heterogeneity of the causal effect (when applicable). Specific details about the simulation scenarios can be found in Appendix D, though we summarize the key points here. We let treatment status at time $t$ depend on $Y_{i,t-1}(0)$, and follow a staggered adoption setting where a unit remains treated once they are treated. We vary the degree of confounding by $Y_{i,t-1}(0)$ into three categories we refer to as no confounding, moderate confounding, and high confounding. We set a small average treatment effect and vary the degree of heterogeneity of this effect across units, split into categories denoted by no heterogeneity, moderate heterogeneity, and high heterogeneity. The moderate and high heterogeneity settings represent scenarios where the probability that a unit has an individual treatment effect either twice as large, or the opposite sign, as the average treatment effect is set to be 0.1 and 0.33, respectively. It is also important to emphasize that the only assumption placed on simulated data that coincides with our modeling assumptions is that the effects are homogeneous with respect to time. Every other aspect of the simulated data comes from the real, observational data, and may or may not fit well into the autoregressive framework we presented.
5.1 Target estimand and competing estimators
We explore four different estimators in this manuscript: 1) The autoregressive model that assumes a linear autoregressive model on the rates of the outcome that incorporates two time period lags, 2) The two way fixed effects estimator, 3) the synthetic control estimator for staggered adoption developed in Ben-Michael, Feller and Rothstein (2022), and 4) the difference in difference estimator for staggered adoption developed by Callaway and Sant’Anna (2021). Note that we use the linear autoregressive model here that uses a normal approximation to the rates of the outcome instead of directly modeling counts as described in Section 3.4. In Appendix D we run the same simulations using negative binomial autoregressive models and find largely similar results. Further, to ensure a fair comparison across different estimation procedures, our target of inference throughout is the sample treatment effect on the treated:
$\frac{1}{n_{t}}\sum_{i\in\mathcal{N}_{t}}\sum_{m=0}^{2}\bigg{\{}Y_{i,T_{i}+m}(1)-Y_{i,T_{i}+m}(0)\bigg{\}}.$
This estimand targets the average treatment effect on the treated units in the first three time periods after policy adoption. For more details on how to perform inference on sample level estimands within our framework, see Appendix C or read Ding and Li (2018) for a general discussion. Most importantly, however, this estimand is easily estimated using any of the synthetic control, DiD, and autoregressive models we consider. Interpretation of the TWFE estimator may be complicated in this setting (Imai and Kim, 2021), but we include it due to its widespread use. We calculate standardized mean squared error (MSE), which in any simulation scenario is the MSE of an estimator divided by the minimum MSE across all estimators, leading to the optimal estimator having an MSE of 1. We also calculate 95% interval coverage rates and the power to detect a nonzero effect, which is calculated as the proportion of 95% intervals for which zero is not included and the interval correctly estimates the sign of the treatment effect.
@@ 5.2 Results on simulation using opioid prescribing data
Figures 3 - 6 display the results. The best performing estimators are the autoregressive model and synthetic control approach, followed by the difference in differences estimator, and then the TWFE estimator. The 95% interval coverage is good for the autoregressive, DiD, and SC approaches with the autoregressive approach only falling below 95% coverage with high heterogeneity or large amounts of confounding. The autoregressive approach has the best MSE in all scenarios with a standardized MSE of 1, and it is substantially more efficient in certain scenarios. This highlights the bias-variance trade-off inherent to using these models as they can be significantly more efficient than existing estimators that are not model based. This boost in efficiency is most notable in Figure 6, where we display power. Both the DiD and SC approaches have effectively no power to detect the nonzero treatment effects, while the autoregressive approach detects this effect in over half of the simulated data sets. This is crucial, because policy effects are typically very small in state-level opioid policies and it appears that existing approaches are not well-suited to detecting these effects. Another key takeaway from the simulation is that reasonable amounts of heterogeneity does not appear to negatively impact the performance of the autoregressive model, consistent with the results of Section 2.3 suggesting that only very large amounts of heterogeneity will lead to poor performance of these models.
@@ 6 Examining the impact of state policies on opioid prescriptions
We now examine the impact of four state-level policies on the number of opioid prescriptions per capita in each state over the years 2006-2021. We first provide background on the data sources used and the questions we set out to answer, and then utilize the proposed autoregressive model to estimate the effects of these policies.
6.1 Covariates, policies, and outcome of interest
To measure the key policies of interest, we utilized OPTIC-Vetted Policy Data Sets from the RAND-USC Opioid Policy Tools and Information Center (OPTIC) to provide us with high quality policy measures (RAND-USC Schaeffer Opioid Policy Tools and Information Center, 2024a, b, c). We choose to focus on four different policies that have been examined in prior research but are rarely considered jointly. Naloxone is a potential life saving treatment for opioid overdoses that works by reversing the effects of opioids (Cawley and Dragone, 2023; Chamberlain and Klein, 1994), though individuals face systemic stigma or lack access to traditional healthcare systems, making it difficult to obtain naloxone (Green et al., 2020; Smart and Grant, 2021). To address these barriers, every state has adopted some type of NAL by 2018 (Smart, Pardo and Davis, 2021). There are two broad classes of naloxone access laws (NALs Smart et al. (2023)): those which allow distribution through a standing or protocol order or those which allow distribution through a pharmacist prescriptive authority (Davis and Carr, 2015, 2017).
A third policy we study is prescription drug monitoring programs (PDMPs), which are electronic databases that track prescribed and dispensed controlled substances within a state (Schuler et al., 2020). By giving healthcare providers information about a patient’s prescription drug history, PDMPs may help identify patients who are at risk for opioid misuse or overdose, potentially increasing early intervention and support (Horwitz et al., 2018). Based on the literature on PDMPs, state laws that specify that prescribers must access the PDMP before writing a prescription to a new/existing patient tends to allow for more effective forms of PDMPs (Pacula and Stein, 2020; RAND-USC Schaeffer Opioid Policy Tools and Information Center, 2024b). Thus, we define our PDMP laws as those which have this must access component. Lastly, many states have legalized marijuana for medical purposes (Pacula et al., 2015). Medical marijuana laws have the potential to reduce opioid use and opioid-related harms by providing an alternative treatment option for chronic pain and other conditions that are commonly treated with opioids (Powell, Pacula and Jacobson, 2018). For all of these policies, we record the fraction of the year each state has such a law in place. For a summary of the prevalance of these laws, and how this varies over time, see Figure 7
Figure 7: Policy wheels showing the timeline of policy adoption
We examined the impact of these policies on annual CDC opioid prescribing rates per a population of 100,000 people from 2006-2021, constructed from the QuintilesIMS Transactional Data Warehouse (TDW). These data represent approximately 59,000 retail (non-hospital) pharmacies, which dispense nearly 88% of all retail prescriptions in the U.S. Opioid prescriptions were identified using the National Drug Code, while cough and cold formulations containing opioids and buprenorphine, an opioid partial agonist used for treatment of opioid use disorder as well as for pain, were not included. For rates, annual resident population denominator estimates were obtained from the Population Estimates Program, U.S. Census Bureau. An illustration of opioid prescription rates over time by state can be found in Appendix E. Finally, we included seven time-varying covariates in our models that include unemployment rates, median income, and poverty rates, as well as percentages of non-Hispanic white individuals, females, individuals under 20 years old, and individuals greater than 60 years old. These covariates were selected to capture a range of sociodemographic factors that might confound the effects of our policies on the outcomes.
@@ 6.2 Single policy analysis
First, we analyze the impact of each policy on prescribing separately, which facilitates comparison with existing approaches designed for one policy only. Additionally, our policy variables are proportions between 0 and 1, but for this analysis, we dichotomize policy variables based on whether the state has any law in place that year.
@@ 6.2.1 Model formulation and estimands of interest
Our estimand of interest is the sample average treatment effect on the treated used in Section 5, which represents the average impact of the policy on the treated units in the first 3 years it is enacted. We utilize four distinct estimators of this quantity: 1) a DiD estimator allowing for staggered adoption, 2) the augmented synthetic control approach that allows for staggered adoption, 3) the autoregressive model incorporating time-varying covariates, and 4) the autoregressive model without time-varying covariates. For the autoregressive models, we must specify the specific form of our model as well as the prior distributions to be used. We assume that $k=2$ and that a second order autoregressive model is sufficient, and we enforce that only the policy indicator at the most recent three time points affects the outcome. The autoregressive coefficients are assigned standard uniform prior distributions, while all regression parameters, including those that govern the treatment effect, are assigned noninformative normal priors centered at zero with a variance of 100, and the variance parameters are assigned inverse gamma priors with parameters 0.001 and 0.001.
6.2.2 Results
The results for each of the four policies under each of the four estimators is given in Table 1. Note there is no estimate for the augmented synthetic control estimator for the medical marijuana policy, as many states had already enacted this policy at the start of our study and therefore this policy is not amenable to synthetic controls. For both NAL policies as well as medical marijuana, there is a fairly consistent story across all estimators, which is that they do not have an impact on opioid prescriptions. All estimators estimate a negative effect of PDMP laws on prescriptions suggesting must access PDMP laws decrease opioid prescribing, with both the DiD and autoregressive models showing a statistically significant effect. In general across all four policies, the autoregressive models tend to have much smaller interval widths, highlighting the efficiency gains that come with this approach.
In addition to the sample average treatment effect that is averaged over the first three time periods, we also estimate a time specific sample average treatment effect that does not average over time periods. We show estimates of this quantity for the autoregressive model with covariates in Figure 8. We see no effect at any time period for NAL or medical marijuana laws, which is expected given the previous results. We see that the effect of PDMP laws becomes more pronounced after the first year. This could be due to the fact that we dichotomized the treatment variable, but in the first year in which it is enacted, it is commonly only enacted for a fraction of the year.
@@ 6.3 Examining multiple policies simultaneously
We can also use the proposed approach to estimate the effect of multiple policies simultaneously, allowing us to see whether certain policies are more (or less) effective if implemented in tandem with other policies. We no longer dichotomize treatment variables and instead use the proportion of that year that the policy was implemented as our main treatment variables. We use a similar model formulation as for the single policy case as we utilize $k=2$ autoregressive lags, though here we only allow the treatments at the current time point to affect outcome. This is done partially to reduce the complexity of the model in order to allow us to study multiple policies and their interactions simultaneously. For this analysis, we focus on two policies of interest: must access PDMPs and medical marijuana laws. Studying all four laws simultaneously is exceedingly difficult with only 51 units and leads to serious issues with positivity violations as there are combinations of the four policies for which we have no data. We focus on the must access PDMP and medical marijuana laws specifically, because they are the two most likely to affect opioid prescription rates. Naloxone access laws are primarily expected to reduce opioid mortality by intervening in cases of an overdose, rather than by reducing prescriptions.
The estimand of interest is similar to the one studied for single policy analyses, but now potential outcomes are denoted by $Y_{it}(a,m)$, which is the outcome we would observe for unit $i$ at time $t$ had they been exposed to levels of the two policies given by $a$ and $m$. We look at a sample average potential outcome given by $SAPO(a,m)=\frac{1}{n}\sum_{i=1}^{n}Y_{i,T_{i}}(a,m)$, for all combinations of $a$ and $m$. This shows the average outcome, across all states, for each combination of the two policies. Moving forward, the first policy will represent must access PDMP laws, while the second policy will represent medical marijuana laws.
@@ 6.3.1 Results
The estimates and corresponding 95% credible intervals for SAPO($a$, $m$) for the four combinations of policies is given in Figure 9. The values of SAPO(00, 00) and SAPO(00, $1$) are both high indicating that medical marijuana does little to reduce opioid prescriptions. We also see that SAPO($1$, 00) is much lower, which re-iterates what was seen in the single policy analysis, which is that must access PDMP laws alone are able to substantially reduce opioid prescriptions at the state level. Interestingly, the value of SAPO($1$, $1$) is moderately higher than SAPO($1$, 00), which indicates that the presence of medical marijuana policies can dampen the influence of must access PDMPs. To see this, we can also examine the posterior distribution of the coefficients corresponding to the policy variables in the autoregressive model. The treatment effect component of the model is given by $A\theta_{pdmp}+M\theta_{mm}+(A\times M)\theta_{int}$, and the posterior distribution of these three parameters can be seen in Appendix E. The interaction term has posterior density almost entirely greater than 0 (95% CI: 0.50, 4.89), highlighting that medical marijuana laws do temper the effect of the must access PDMP laws on opioid prescriptions.

### v2

@@ 6 Simulation studies
We evaluate the performance of the proposed approach against state-of-the-art methods for panel data causal inference in a simulation study. We explore four different estimators in this manuscript: 1) The linear autoregressive model described in Sections 3 and 4, 2) The two way fixed effects estimator, 3) the synthetic control estimator for staggered adoption developed in Ben-Michael, Feller and Rothstein (2022), and 4) the difference in differences estimator for staggered adoption developed by Callaway and Sant’Anna (2021). In Section 6.1, we compare these approaches on simulations using the observed opioid data. Then, in 6.2 we use fully simulated data generated under linear factor models to evaluate the performance of our estimator in the presence of time-varying unmeasured confounding.
@@ 6.1 Results on simulation using opioid prescribing data
Instead of simulating outcomes from a data generating process we fully specify, we first utilize existing outcome data from our application to ensure that the results reflect performance in realistic data and are more relevant to our setting. Specifically, we let our observed outcome data represent $Y_{it}(0)$, and then we simulate treatment effects to generate $Y_{it}(1)$. Our data consist of 51 units and 16 time periods, and we choose $n_{t}=20$ units to be treated in all simulations. Our observed outcomes are the number of opioids prescribed by each state (and Washington DC) and year. We do not vary the outcome values across simulations, but instead vary which units are treated, when they are treated, and the heterogeneity of the causal effect (when applicable). Specific details about the simulation scenarios can be found in Appendix D, though we summarize the key points here. We let treatment status at time $t$ depend on $Y_{i,t-1}(0)$, and simulate a staggered adoption setting where a unit remains treated once they are treated. We vary the degree of confounding by $Y_{i,t-1}(0)$ into three categories we refer to as no confounding, moderate confounding, and high confounding. We set a small average treatment effect and vary the degree of heterogeneity of this effect across units, split into categories denoted by no heterogeneity, moderate heterogeneity, and high heterogeneity. The moderate and high heterogeneity settings represent scenarios where the probability that a unit has an individual treatment effect either twice as large, or the opposite sign, as the average treatment effect is set to be 0.1 and 0.33, respectively. It is also important to emphasize that we do not simulate the outcomes and therefore they may or may not fit well into the autoregressive framework we presented. Throughout this section, we adjust for the prior two time periods’ counterfactual outcomes. We also use the linear autoregressive model here that uses a normal approximation to the rates of the outcome instead of directly modeling counts as described in Section 4.5. In Appendix D we run the same simulations using negative binomial autoregressive models and find largely similar results.
To ensure a fair comparison across different estimation procedures, our target of inference throughout is the sample treatment effect on the treated described in Section 3 – this estimand is easily estimated using any of the synthetic control, DiD, and autoregressive models we consider. Interpretation of the two-way fixed effects (TWFE) estimator may be complicated in this setting (Imai and Kim, 2021), but we include it due to its widespread use. We calculate relative mean squared error (MSE), which in any simulation scenario is the MSE of an estimator divided by the minimum MSE across all estimators, leading to the optimal estimator having an MSE of 1. We also calculate 95% interval coverage rates and the power to detect a nonzero effect, which is calculated as the proportion of 95% intervals for which zero is not included and the interval correctly estimates the sign of the treatment effect.
@@
Figures 4 - 7 display the results. The best performing estimators are the autoregressive model and synthetic control approach, followed by the difference in differences estimator, and then the TWFE estimator. The 95% interval coverage is good for the autoregressive, DiD, and SC approaches with the autoregressive approach only falling below 95% coverage with high heterogeneity or large amounts of confounding. The autoregressive approach has the best MSE in all scenarios with a relative MSE of 1, and it is substantially more efficient in certain scenarios. This highlights the bias-variance trade-off inherent to using these models as they may be significantly more efficient than existing estimators that are not model based, although they may be more sensitive to model misspecification. This boost in efficiency is most notable in Figure 7, which displays power. Both the DiD and SC approaches have effectively no power to detect the nonzero treatment effects, while the autoregressive approach detects this effect in over half of the simulated data sets. This is crucial, because policy effects are typically very small in state-level opioid policies and it appears that existing approaches are not well-suited to detecting these effects. Another key takeaway from the simulation is that reasonable amounts of heterogeneity do not appear to negatively impact the performance of the autoregressive model, consistent with the results of Section 4.1.
@@
6.2 Results using fully simulated data
We also present results using fully simulated data to highlight the performance of our approach under the linear factor model that was described in Section 5.1, and that is commonly used to justify synthetic control approaches, and that nests difference-in-differences as a special case. Specifically, we generate data from a factor model given by
$Y_{it}(0)=\bm{\lambda}_{t}\bm{\mu}_{i}+\epsilon_{it},$
where we generate the latent factors $\bm{\lambda}_{t}$ from independent standard normal distributions. The time-varying coefficients $\bm{\lambda}_{t}$ are drawn randomly from a standard uniform distribution, and $\epsilon_{it}\stackrel{{\scriptstyle iid}}{{\sim}}\mathcal{N}(0,1)$. The number of latent factors is given by $r$, which we vary in $\{2,5,10\}$ to show the impact of increasing numbers of latent factors. Throughout, we set the number of units to be $n=20$, which are measured over $T=12$ time points. We explore two distinct simulation scenarios using the factor model above, which vary in the treatment assignment mechanism. Note that treatment status for one unit in this scenario can switch between the treatment and control condition over time unlike in the previous simulation setting. Letting $g(\cdot)$ be the standard logistic link function, these two propensity score models are given below.
  1. 1.
    A constant assignment model with probabilities given by
    $P(A_{it}=1\mid\bm{\mu}_{i})=g\bigg(-0.5+\sum_{j}\mu_{ij}\bigg)$
  2. 2.
    A time-varying assignment model where $\bm{\xi}_{t}$ are randomly assigned from independent, standard uniform distributions, and the propensity score is given by
    $P(A_{it}=1\mid\bm{\mu}_{i})=g\bigg(-0.5+\bm{\xi}_{t}^{\top}\bm{\mu}_{i}\bigg)$
We fit our proposed approach and vary the number of lags that we adjust for in $\{2,4,6,8,10\}$. The point estimates across 500 replications can be found for each scenario in Figure 8.
Our approach is able to obtain approximately unbiased results in the constant propensity score setting, but is unable to achieve these same results in the time-varying propensity score setting. Moreover, the bias in the constant propensity score setting decreases with the number of lags included in the model, although we increasingly lose efficiency, reflected by the increasing spread of the point estimates. The bias results are consistent with our results in Proposition 5 in Appendix G and with (Abadie, Diamond and Hainmueller, 2010). Heuristically, the failure of our model in the time-varying setting occurs because the distribution of factors across the treatment and control groups change over time; as a result, controlling for the prior counterfactual outcomes cannot simultaneously balance all unobserved factors across the treatment and comparison groups in each time period. [^9] We expect that this failure would occur in any model that does not either implicitly or explicitly allow the treatment assignment model to vary over time. By contrast, in the constant propensity score setting, this distribution remains constant over time, and the lagged outcomes appropriately proxy for the unobserved factors. Moreover, the more lags we control for, the better our models balance these factors appropriately across the treatment and comparison groups.
Figure 8: Distribution of point estimates for simulations from latent factor models.
@@ 7 Examining the impact of state policies on opioid prescriptions
We now estimate the effect of the four state-level policies described in Section 2 on the number of opioid prescriptions per capita in each state over the years 2006-2017. We first focus on estimating the effects of single policies and compare effect estimates with existing approaches to inference in this problem. We then estimate the effect of multiple policies simultaneously using the proposed autoregressive framework.
@@ 7.1 Single policy analysis
First, we analyze the impact of each policy on prescribing separately, which facilitates comparison with existing approaches designed for one policy only. Additionally, we have information about the proportion of each year that the policy is implemented for each state, but for this analysis, we dichotomize policy variables based on whether the state has any law in place that year. This allows for implementation of existing methods designed for binary treatments, but it is worth noting that our approach applies even when $A_{it}$ is a proportion. We explore this further in the multiple policy analysis in Section 7.2.
@@ 7.1.1 Model formulation and estimands of interest
We estimate the sample average treatment effect on the treated, which represents the average impact of the policy on the treated units in the first 3 years it is enacted. We utilize four distinct estimators of this quantity: 1) a DiD estimator allowing for staggered adoption, 2) the augmented synthetic control approach that allows for staggered adoption, 3) the autoregressive model incorporating time-varying covariates, and 4) the autoregressive model without time-varying covariates. For the autoregressive models, we must specify the specific form of our model as well as the prior distributions to be used. We assume that $k=2$ and that a second order autoregressive model is sufficient, and we enforce that only the policy indicator at the most recent three time points affects the outcome. The autoregressive coefficients are assigned standard uniform prior distributions, while all regression parameters, including those that govern the treatment effect, are assigned noninformative normal priors centered at zero with a variance of 100, and the variance parameters are assigned inverse gamma priors with parameters 0.001 and 0.001.
7.1.2 Estimates of average causal effects
The results for each of the four policies under each of the four estimators is given in Table 3. Note there is no estimate for the augmented synthetic control estimator for the medical marijuana policy, as many states had already enacted this policy at the start of our study and therefore synthetic controls cannot estimate this policy effect. For both NAL policies as well as medical marijuana, the results are fairly consistent, showing that they do not have an impact on opioid prescriptions. All estimators estimate a negative effect of PDMP laws on prescriptions suggesting must access PDMP laws decrease opioid prescribing, with both the DiD and autoregressive models showing a statistically significant effect. In general across all four policies, the autoregressive models tend to have much smaller interval widths, highlighting the efficiency gains of this approach.
In addition to the sample average treatment effect, we also estimate a time specific sample average treatment effect that does not average over time periods. Figure 9 shows estimates of this quantity using the autoregressive model, revealing no effect at any time period for NAL or medical marijuana laws. The effect of PDMP laws, however, becomes more pronounced after the first year. This may be due to the fact that we dichotomized the treatment variable, but in the first year in which it is enacted, it is commonly only enacted for a fraction of the year.
@@ 7.1.2 Estimates of average causal effects
7.1.3 Exploring treatment effect heterogeneity
We also explore effect heterogeneity, using the methodology described in Section 4.1 and let $V_{it}=X_{it}$ so that we study heterogeneity across all observed covariates. First, we examine whether accounting for heterogeneity leads to different average causal effect estimates than those in Section 7.1.2. The estimates are largely similar. The average effects and corresponding credible intervals are for each policy are: 1) NAL protocol standing: 0.36 (-0.69, 1.47); 2) NAL prescriptive authority: -0.26 (-2.31, 1.78); 3) Must Access PDMP: -2.79 (-4.72, 1.02); and 4) Medical marijuana: -0.03 (-1.14, 1.10). While these yield the same substantive conclusions as before, the widths of the 95% credible intervals for all four policies are wider, illustrating the trade-off from the additional model complexity.
We also examine the extent of heterogeneity for any of the policies by examining the coefficients $\nu_{j}$, which highlight how the treatment effect function changes with each covariate. We show these coefficients for the Must Access PDMP policy, which is the only policy to show a marginal effect on opioid rates. Figure 10 displays the results. All intervals, whether pointwise or simultaneous, contain zero, indicating that we cannot reject the null hypothesis of no treatment effect heterogeneity. The interval for percentage in poverty is mostly negative, indicating that states with higher numbers of people in poverty would have more pronounced, negative effects of must access PDMP policies.
Figure 10: 95% credible intervals for the coefficients of the treatment effect heterogeneity function when examining Must Access PDMP policies.
@@ 7.2 Examining multiple policies simultaneously
We may also use the proposed approach to estimate the effect of multiple policies simultaneously, allowing us to see whether certain policies are more (or less) effective if implemented in tandem with other policies. We no longer dichotomize treatment variables and instead use the proportion of that year that the policy was implemented as our main treatment variables. We use a similar model formulation as for the single policy case as we utilize $k=2$ autoregressive lags, though we now only allow the treatments at the current time point to affect outcome. We do this partially to reduce the complexity of the model to allow us to study multiple policies and their interactions simultaneously. For this analysis, we focus on two policies of interest: must access PDMPs and medical marijuana laws. We focus on the must access PDMP and medical marijuana laws specifically, because they are the two most likely to affect opioid prescription rates. Naloxone access laws are primarily expected to reduce opioid mortality by intervening in cases of an overdose, rather than by reducing prescriptions. [^10] Studying all four laws simultaneously is exceedingly difficult with only 51 units and leads to serious issues with positivity violations as there are combinations of the four policies for which we have no data.
The estimand of interest is similar to the one studied for single policy analyses, but now potential outcomes are denoted by $Y_{it}(a,m)$, which is the outcome we would observe for unit $i$ at time $t$ had they been exposed to levels of the two policies given by $a$ and $m$. We look at a sample average potential outcome given by $SAPO(a,m)=\frac{1}{n}\sum_{i=1}^{n}Y_{it}(a,m)$, for all combinations of $a$ and $m$. We focus on $t=6$, which corresponds to the year 2011, though similar results are obtained examining different years. This shows the average outcome, across all states, for each combination of the two policies. Moving forward, the first policy will represent must access PDMP laws, while the second policy will represent medical marijuana laws.
@@ 7.2.1 Results
The estimates and corresponding 95% credible intervals for SAPO($a$, $m$) for the four combinations of policies is given in Figure 11. The values of SAPO(0, 0) and SAPO(0, 11) are both high, indicating that medical marijuana does little to reduce opioid prescriptions. SAPO($1$, 0) is much lower, which re-iterates what was seen in the single policy analysis, which is that must access PDMP laws alone are able to substantially reduce opioid prescriptions at the state level. Interestingly, the value of SAPO($1$, 11) is moderately higher than SAPO($1$, 0), suggesting that the presence of medical marijuana policies can dampen the influence of must access PDMPs. To see this, we may examine the posterior distribution of the coefficients corresponding to the policy variables in the autoregressive model. The treatment effect component of the model is given by $A\theta_{pdmp}+M\theta_{mm}+(A\times M)\theta_{int}$, and the posterior distribution of these three parameters can be seen in Appendix E. The interaction term has posterior density almost entirely greater than 0 (95% CI: 0.50, 4.89), highlighting that medical marijuana laws temper the effect of the must access PDMP laws on opioid prescriptions.